OpenStudy (anonymous):

What is f(–3) if f(x) = |2x + 1| + 5? can someone help with this .i'll give medal :)

OpenStudy (solomonzelman):

you just need to plug in -3 instead of x into the function.

OpenStudy (solomonzelman):

\(\large\color{black}{ f(x) = \left| 2x + 1 \right| + 5 }\) For example, if I wanted to find, \(\large\color{black}{ f(-5) }\), this is what I do: \(\large\color{black}{ f(\color{red}{-5}) = \left| 2(\color{red}{-5}) + 1 \right| + 5 }\) \(\large\color{black}{ f(\color{red}{-5}) = \left| -10+ 1 \right| + 5 }\) \(\large\color{black}{ f(\color{red}{-5}) = \left| -9 \right| + 5 }\) then you know that for any positive number \(\large\color{black}{ a }\), \(\large\color{black}{ \left| -a \right| =a }\) So, you get, \(\large\color{black}{ f(\color{red}{-5}) = 9 + 5 }\) \(\large\color{blue}{ f(\color{red}{-5}) = 14 }\)

OpenStudy (anonymous):

oh ok thank you

OpenStudy (solomonzelman):

Anytime!